Find the Taylor polynomial of order 4 (at x = 0) for the function : f(x) = ex^2



1
Expert's answer
2021-07-30T14:35:14-0400


In general,

the Taylor polynomial of order n for f(x) centered at x=a is given by

"T_n(x)=f(a)+f'(a)(x-a)+\frac{f''(a)}{2}(x-a)^2+\frac{f'''(a)}{6}(x-a)^3...+\frac{f^{(n-1)}(a)}{(n-1)!}(x-a)^{n-1}+\frac{f^{(n)}(a)}{n!}(x-a)^n"

Therefore,

Find the respective derivatives of f(x)

f(x)="e^{x^2}" ;f(0)=1

f'(x)="2xe^{x^2}" ;f'(0)=0

f''(x)="(4x^2+2)e^{x^2}" ;f''(0)=2

f'''(x)="(8x^3+12x)e^{x^2}" ;f'''(0)=0

f''''(x)="(16x^4+48x^2+12)e^{x^2}" ;f''''(0)=12

Hence,

T4(x)="1+0+x^2+0+\frac12x^4"

"T_4(x)=1+x^2+\frac12x^4"